A machine gun fires a bullet of mass 40 g with a velocity 1200 m/sย  The man holding it, can exert .ย  a maximum force of 144 N on the gun. How many bullets can he fire per second at the most?

Grade:11

2 Answers

upender surepally
11 years ago

The force exerted by machine gun on man''s hand in firing a bullet = change in momentum per second on a bullet or rate of change of momentum =(40/1000*)1200= 48 N .
The force exerted by man on machine gun = 144 N
Hence, number of bullets fired =  3

Shivam Bhagat
11 years ago

According to newtons 2nd law..... F=m(v-u)/t

144=0.04*n*1200/1           (per second...so t=1s)

n=3 bullets / sec

ALL THE BEST..!